# Learning What Reinforcement Learning Can’t: Interleaved Online Fine-Tuning for hardest Questions

### v1

@@ 3.5 Extension to More Models
Qwen-Math-1.5B | 3.0 | 4105 | 1.4 | 4291 | 19.4 | 3480 | 44.4 | 2637 | 16.7 | 3487 | 5.0 | 6260 | 15.0 | 4043 |
Qwen-Math-1.5B-Instruct | 8.1 | 5722 | 6.6 | 4468 | 36.9 | 4614 | 66.0 | 6148 | 30.7 | 5526 | 24.2 | 8360 | 28.8 | 5806 |
SFT | 12.7 | 18060 | 13.0 | 19271 | 40.9 | 12538 | 71.8 | 5879 | 33.8 | 12906 | 23.5 | 11767 | 32.6 | 13403 |
RL | 10.2 | 1487 | 8.7 | 1187 | 44.7 | 796 | 71.4 | 423 | 34.5 | 832 | 28.7 | 756 | 33.0 | 913 |
ReLIFT | 13.1 | 6294 | 8.8 | 5866 | 42.5 | 3856 | 73.6 | 2750 | 36.0 | 3937 | 30.1 | 3969 | 34.0 | 4421 |
Qwen-7B | 5.1 | 3146 | 1.4 | 3518 | 22.6 | 2611 | 37.2 | 1811 | 19.0 | 2429 | 39.8 | 434 | 20.9 | 2324 |
Qwen-7B-Instruct | 10.3 | 1299 | 6.5 | 562 | 34.9 | 1043 | 67.6 | 605 | 29.8 | 907 | 53.3 | 7413 | 33.7 | 1804 |
SFT | 15.7 | 19363 | 18.6 | 18070 | 49.8 | 11349 | 80.8 | 4453 | 44.3 | 11528 | 54.6 | 5188 | 43.9 | 11658 |
RL | 13.3 | 4371 | 4.2 | 3988 | 40.2 | 2845 | 65.4 | 1971 | 31.4 | 2852 | 56.6 | 180 | 35.2 | 2867 |
ReLIFT | 19.1 | 12197 | 14.7 | 9284 | 51.9 | 5909 | 81.6 | 2395 | 46.1 | 6160 | 56.4 | 2550 | 45.0 | 6415 |
To evaluate the general applicability of our method, we extend ReLIFT to both a smaller base model, Qwen2.5-Math-1.5B Yang et al. (2024), and a weaker base model, Qwen2.5-7B Team (2024). Table 3 presents the performance across five challenging competition-level math benchmarks and one out-of-domain benchmark. On both models, ReLIFT consistently achieves substantial improvements, surpassing both SFT and RL. Across all benchmarks, our method ranks first or second, while maintaining an acceptable response length. These results underscore the applicability of ReLIFT to achieve both higher performance and better generalization.
4 Related Works
4.1 RL-Driven Emergence of Reasoning Abilities
Reinforcement learning has recently emerged as a promising approach for enhancing the reasoning capabilities of large language models (LLMs). Early breakthroughs, such as O1 Jaech et al. (2024), DeepSeek R1 Guo et al. (2025), and Kimi 1.5 Team et al. (2025), have demonstrated that applying Reinforcement Learning with Verifiable Rewards (RLVR) can lead to significant improvements in complex reasoning tasks. Notably, DeepSeek R1 further shows that applying RLVR directly to a base model—without the need for intermediate supervised fine-tuning—can also effectively enhance the reasoning abilities of LLMs.
Building on these foundations, Light-r1 Xie et al. (2025) incorporated curriculum learning to further refine reasoning skills, while Logic-rl Xie et al. (2025) adopted logic-driven reward functions to improve general reasoning ability. Deepscaler Luo et al. (2025) introduced Iterative Context Lengthening, a method that trains models with progressively longer contexts as their performance improves, thereby reducing both monetary costs and end-to-end training time. In addition, Oat-Zero Liu et al. (2025) was trained from Qwen2.5-Math-7B using rule-based rewards, and proposed removing the standard deviation in GRPO advantage computation as well as token-level normalization in policy loss computation. PRIME-Zero Cui et al. (2025) leveraged policy rollouts and outcome labels, utilizing implicit process rewards to further enhance model performance.
4.2 Training Paradigms
Despite significant advances in reinforcement learning (RL), Yue et al. Yue et al. (2025) argue that RL with value-based rewards (RLVR) does not enable large language models (LLMs) to acquire genuinely new reasoning abilities beyond those present in their base models. Instead, RL primarily enhances performance by biasing the model toward reasoning paths it already knows that are more likely to yield rewards. In other words, most RL methods optimize within the model’s existing knowledge rather than facilitating the acquisition of new information, thereby limiting the LLM’s overall capabilities.
In contrast, directly fine-tuning on high-quality demonstration data can introduce new knowledge and abilities to the model, and this approach is widely used to adapt LLMs to downstream reasoning tasks. However, supervised fine-tuning (SFT) requires high-quality, detailed demonstration data and tends to encourage memorization of the training data, often resulting in poor generalization to out-of-distribution scenarios Chu et al. (2025).
Some approaches attempt to combine supervised fine-tuning (SFT) and reinforcement learning (RL). Typically, these methods first use SFT to stabilize the model’s output format, then apply RL to further improve performance and potentially unlock more advanced reasoning abilities in LLMs Guo et al. (2025); Wen et al. (2025). However, such combined methods remain limited: SFT is highly dependent on the quality of demonstration data, and the subsequent RL phase is still confined to the knowledge already present in the SFT-trained model, making it difficult for the model to acquire genuinely new information.
LUFFY Yan et al. (2025) seeks to address these limitations by augmenting RL with off-policy reasoning traces, dynamically balancing imitation and exploration through a combination of off-policy demonstrations and on-policy rollouts during training. However, this approach requires the preparation of off-policy data in advance and relies heavily on probability-sharpening techniques, which may restrict its applicability. In contrast, our work proposes a simpler and more direct method that combines RL and SFT to leverage the strengths of both paradigms. Our approach encourages LLMs to explore when possible and to incorporate new information as needed, thereby pushing RL-driven reasoning beyond the boundaries of existing knowledge.
5 Conclusion
In summary, we systematically analyze the respective strengths of RL and SFT for LLM reasoning, finding that RL is more effective for learning easier questions while SFT is crucial for solving challenging questions. Building on these insights, we introduce ReLIFT, an approach that interleaves RL with online Fine-Tuning on the hardest questions, achieving state-of-the-art results with significantly less demonstration data and more concise solutions. Future work will focus on scaling ReLIFT to larger models and developing more effective strategies for coordinating SFT and RL to further advance LLM reasoning and generalization.
Appendix A Experiment Details
A.1 Main Experiment Details
Training. In addition to Qwen2.5-Math-7B, we extend ReLIFT to both Qwen2.5-Math-1.5B Yang et al. (2024) and Qwen2.5-7B Team (2024). Since the maximum sequence length for Qwen2.5-Math-1.5B is 4096, which is insufficient for our tasks, we follow the approach used for Qwen2.5-Math-7B by increasing the RoPE theta from 10,000 to 40,000 and expanding the window size to 16,384. To ensure fairness, we maintain 8 rollouts per prompt for all RL-trained models. The learning rate is fixed at . All training experiments are conducted on 2 8 A800 GPUs. We train all RL models for 500 steps and all SFT models for three epochs. For all experiments, we use verl Sheng et al. (2024).
Evaluation. All evaluations are conducted using vllm Kwon et al. (2023), with the temperature set to 0.6 and the maximum number of tokens set to 8192. Results are verified using math-verify Face (2024). For the baselines, we independently reproduce and verify the results of the LUFFY Yan et al. (2025), whereas the results for the other zero-RL models are reported as in the LUFFY paper.
Chat Template. To encourage the base model to reason more effectively, following Yan et al. (2025), we adopt a complex chat template as shown in Figure 5. All training experiments and evaluation utilize the same chat template.
A.2 Motivation Experiment Details
Dataset. We randomly sample a subset of 8,000 examples from the training dataset described in 3.1 as the training set, and select 1,000 examples as the validation set. RL and SFT are both conducted on this subset dataset for 120 steps, using the same parameters as in our main experiments.

### v2

@@ Appendix D Detailed Experiment Results
| ReLIFT | 28.3 | 5062 | 22.9 | 4571 | 65.1 | 3540 | 87.4 | 2241 | 57.8 | 3352 | 53.9 | 2248 | 52.6 | 3502 |
| ReLIFT(all) | 7.8 | 7647 | 9.0 | 7648 | 31.6 | 6653 | 54.2 | 5320 | 26.4 | 6600 | 13.7 | 6590 | 23.8 | 6743 |
| ReLIFT(uniform) | 23.2 | 6113 | 19.3 | 5318 | 62.3 | 4016 | 86.0 | 1971 | 52.4 | 2926 | 51.4 | 1952 | 49.1 | 3716 |
| ReLIFT(random) | 22.1 | 4616 | 13.9 | 4676 | 59.5 | 5271 | 84.2 | 5820 | 44.7 | 5187 | 48.4 | 6020 | 45.5 | 5268 |
The comprehensive results for ReLIFT(all), ReLIFT(uniform), and ReLIFT(random) are presented in Table 4. These findings indicate that merely alternating between RL and fine-tuning is insufficient; both the scheduling and the selection of fine-tuning training data are crucial components for success.
Appendix E Reasoning Behaivers
To analyze the distinct behaviors of the RL, SFT, and ReLIFT models, we first curate a pool of keywords corresponding to three key cognitive actions: summarizing, rethinking, and planning. We then quantified the frequency of these keywords within the AIME25 responses generated by each model. To ensure a fair comparison across varying response lengths, all keyword counts are normalized by the total number of tokens. As shown in Figure 8, our analysis reveals several key differences in how the models approach problem-solving.
Summarizing: The SFT model exhibits a notably higher usage of summary keywords, suggesting it often prioritizes synthesizing and restating information. The ReLIFT model uses slightly more summary keywords than the RL model, but both are far behind SFT.
Rethinking: In contrast, the ReLIFT model uses significantly fewer keywords related to rethinking or re-evaluation. This finding suggests a higher level of efficiency and confidence in its initial approach, requiring less internal correction or revision during the problem-solving process.
Planning: The most striking difference is in the use of planning keywords. The ReLIFT model demonstrates a substantial increase in these keywords, indicating a more robust and sophisticated approach to strategic planning. This suggests that ReLIFT is not just more efficient but also more capable of developing a logical and powerful problem-solving trajectory.
Appendix F Case Study
AIME25 I Problem 3: The first step to solve this problem is to find three possible combinations , and . The RL response fails to include one possible combination, while the SFT response requires excessive analysis to obtain the three combinations. In contrast, the ReLIFT response is more concise and obtains all three combinations.